# A new characterization of (pre)liminary C*-algebras

### v3

@@ 1. Introduction
This paper is a contribution to the study of the structure and classification of separable C*-algebra. We focus on the class of so-called liminary C*-algebras, also known as liminal or CCR. These are precisely the C*-algebras whose irreducible representations on a Hilbert space have the property that their image coincides with the algebra of compact operators. The more generous requirement that the image contains the algebra of compact operators yields the notion of postliminary C*-algebra, also known as postliminal or Type I. Within the class of postliminary C*-algebras, liminary C*-algebras are precisely those whose primitive spectrum endowed with the Jacobson topology has closed points—i.e., every prime ideal is maximal among the class of prime ideals, and in fact within the class of all ideals [27, Corollary 2.6].
The study of liminary and postliminary C*-algebras goes back to the early days of C*-algebra theory. Its forefathers are Kaplansky [33, 32] and Mackey [42, 39, 40, 41], who pioneered the study of liminary and postliminary C*-algebras in the 1950s. Their motivation included the classification problem for C*-algebras and their irreducible representations. The classification problem of irreducible representations of locally compact Hausdorff topological groups can be seen as a particular instance of the one of C*-algebras, by considering the corresponding group C*-algebra. The theory of liminary and postliminary C*-algebras was further developed in the 1960s by Dixmier [13, 14, 12, 16], Effros [19, 20], Fell [23, 25, 24], and Glimm [28] among others.
An upshot of these works is Glimm’s Theorem from [28] characterizing separable postliminary C*-algebras in terms of the corresponding classification problem for irreducible representations: a separable C*-algebra is postliminary if and only if its irreducible representation are concretely classifiable up to unitary equivalence, and the corresponding Borel structure induced on the spectrum of the C*-algebra is standard.
Since then, the study of the representation theory of C*-algebra has divided into two streams, divided by the type $I$ versus non-type $I$ dichotomy. In the non-type $I$ case, the study of the classification problem for irreducible C*-algebras and their “definable spectra” has refined the analysis from the perspective of Borel complexity theory [35, 52, 22, 31]. In the type $I$ case, more stringent notions have been introduced in the attempt to obtain a complete description of the C*-algebras under considerations and their spectra.
@@
As observed in [21], a type $I_{0}$ and, more generally, a type $I_{n}$ C*-algebra is necessarily liminary, and the same holds for arbitrary countable ordinals. It is also remarked in [21] that there exists liminary C*-algebras that are not type $I_{n}$ for any $n<\omega$. We prove that by considering arbitrary countable ordinal one obtains a hierarchy that is complete, and includes all liminary C*-algebras.
@@
We also prove that the hierarchy is strict, namely, for any countable ordinal $\alpha$ there exist liminary C*-algebras that are type $I_{\alpha}$ and not type $I_{\beta}$ for any $\beta<\alpha$. Such C*-algebras can be chosen to be scattered, namely have countable spectrum, and no infinite-dimensional irreducible representations.
@@
We extend the Enders–Shulman characterization to type $I_{\alpha}$ C*-algebras for an arbitrary countable ordinal $\alpha$. This is obtained in terms of the (ordinal-valued) Pedersen rank of an element of a C*-algebra, generalizing the notion of abelian element (corresponding to the rank 0 case) and global rank $n$ (corresponding to the case of finite ordinals). In turn, this yields a corresponding notion of $\alpha$-subhomogeneous C*-algebra for an arbitrary countable ordinal $\alpha$. Whereas $n$-subhomogeneous C*-algebras have all irreducible representations of dimension at most $n$, when $\alpha$ is infinite an $\alpha$-subhomogeneous C*-algebra can have irreducible representations of arbitrary large, albeit finite, dimension. They are defined by requiring that all their elements have Pedersen rank at most $\alpha$.
We define a C*-algebra to be preliminary if it has no infinite-dimensional representations. A preliminary C*-algebra is, in particular, liminary. This class of C*-algebras has received attention recently, in the work of Courtney and Shulman [10]; see also [30]. As in the case of type $I_{\alpha}$ C*-algebras within liminary C*-algebras, we prove that the hierarchy of $\alpha$-subhomogeneous C*-algebras within preliminary C*-algebras is complete.
@@
Again, this hierarchy is strict, as for any countable ordinal $\alpha$ one can find $\alpha$-subhomogeneous C*-algebras that are not $\beta$-subhomogeneous for any $\beta<\alpha$. Furthermore, one can even find such examples to be scattered. In fact, examples of such C*-algebras were previously considered by Lazar and Taylor [37].
@@ Acknowledgments
We are thankful to Jeffrey Bergfalk, Luigi Caputi, Alessandro Codenotti, Eusebio Gardella, Ilja Gogic, Ivan Di Liberti, Aristotelis Panagiotopoulos, Ebrahim Samei, Tatiana Shulman, and Joseph Zielinski for several helpful conversations.
@@ 2. Some C*-algebra notions
In this section we recall some notions from C*-algebra theory, and particularly the notions of liminary and postliminary C*-algebra; see also [48, 17, 7].
@@ 2.1. Spectrum of a C*-algebra
Let $A$ be a separable C*-algebra. The quasi-state space $Q\left(A\right)$ of $A$ is the space of positive linear functionals on $A$ of norm at most $1$, which is a w*-closed subspace of the dual of $A$ [48, Section 3.2.1]. The space $P\left(A\right)$ of non-zero extreme points of $Q\left(A\right)$ is a $G_{\delta}$ subspace of $Q\left(A\right)$, whence Polish when endowed with the subspace topology [48, Proposition 4.3.2]. The elements of $P\left(A\right)$ are called pure states of $A$.
Let $A$ be a separable C*-algebra. The Borel T-structure (T stands for topological) on $\hat{A}$ is the $\sigma$-algebra generated by the open sets in the Jacobson topology, while the Borel M-structure (M stands for Mackey) on $\hat{A}$ is the one induced from the standard Borel structure on $P\left(A\right)$ by the map $P\left(A\right)\rightarrow\hat{A}$ mapping a pure state $\phi$ to the kernel of the corresponding irreducible representation $\pi_{\phi}$ [48, Theorem 4.3.3]. Then we have that the Borel T-structure is weaker than the Borel M-structure, but they coincide (and they are standard) when $A$ is type $I$ [48, Proposition 6.3.2].
Suppose that $x\in A$. Then $x$ defines a lower semi-continuous function $\check{x}:\mathrm{Prim}\left(A\right)\rightarrow\mathbb{R}$ by $\check{x}\left(J\right):=\left\|x+J\right\|_{A/J}$. In fact, the Jacobson topology on $\mathrm{Prim}\left(A\right)$ is the weakest topology that makes $\check{x}$ lower semi-continuous for all $x$ in (some countable dense subset of) $A$; see [48, Section 4.4].
If $A$ and $B$ are separable C*-algebras, and at least one between $A$ and $B$ is type $I$, then there exists a canonical continuous map from $\hat{A}\times\hat{B}$ to the spectrum of $A\otimes B$ [7, IV.3.4.22]. This is in fact a homeomorphism [7, IV.3.4.28], and it induces a homeomorphism
@@
In particular, $\mathrm{Prim}\left(A\right)$ and $\mathrm{Prim}\left(A\otimes K\left(H\right)\right)$ are homeomorphic, since $\mathrm{Prim}\left(K\left(H\right)\right)$ is a singleton [7, IV.1.2.2].
@@ 2.2. Liminary and postliminary C*-algebras
Let $A$ be a separable C*-algebra, and $\pi$ be an irreducible representation of $A$ on a Hilbert space $H$. Then $\pi$ is CCR if $\pi\left(A\right)\subseteq K\left(H\right)$ (which implies $\pi\left(A\right)=K\left(H\right)$), and GCR if $\pi\left(A\right)\cap K\left(H\right)\neq 0$ (and hence $K\left(H\right)\subseteq\pi\left(A\right)$); see [7, IV.1.3.1]. A C*-algebra is CCR (respectively, GCR) if every irreducible representation of $A$ is CCR (respectively, GCR). A C*-algebra is liminary if and only if it is CCR; see [7, IV.1.3.1]. Recall also that an element $x$ of $A$ is abelian if the hereditary subalgebra $\overline{xAx^{\ast}}$ it generates is commutative [7, IV.1.1.1]. A Fell C*-algebra $A$ (also called type $I_{0}$ C*-algebra) is the C*-subalgebra generated by the abelian elements is equal to $A$ [7, IV.1.1.6]. Every Fell C*-algebra is liminary [7, IV.1.3.2]. A separable C*-algebra is elementary if it is isomorphic to $K\left(H\right)$ for some separable Hilbert space $H$ [7, IV.1.2.1]. Liminary C*-algebras were originally introduced by Kaplansky in [32]; see [48, Section 6.2.13]. Fell C*-algebras were introduced by Pedersen [48, Section 6.1.14], and can be characterized in terms of the so-called Fell condition [7, IV.1.4.17].
@@
Let $A$ be a separable C*-algebra and $\mathcal{C}$ be a class of C*-algebras. A decomposition series with subquotients in $\mathcal{C}$ for $A$ is a chain $\left(R_{\alpha}\right)_{\alpha<\omega_{1}}$ of closed ideals of $A$ such that:
@@
If $A$ is a separable C*-algebra, then it has a largest liminary closed ideal (i.e., a largest element in the collection of closed ideals that are liminal as C*-algebras); see [7, IV.1.3.9]. This can be explicitly defined as the ideal $J_{1}\left(A\right)$ comprising the $x\in A$ such that for every irreducible representation $\pi$ of $A$, $\pi\left(x\right)$ is a compact operator. Then one can recursively define a chain $\left(J_{\alpha}\left(A\right)\right)_{\alpha\leq\omega_{1}}$ of closed ideals, continuous at limits, such that
@@
for $\alpha\leq\beta$. The C*-algebra $A$ is postliminary if $A=J_{\alpha}\left(A\right)$ for some $\alpha<\omega_{1}$; see [7, IV.1.3.9].
A separable C*-algebra also has a largest Fell ideal [7, IV.1.1.8]. This can be explicitly defined as the C*-subalgebra $I_{1}\left(A\right)$ of $A$ generated by its abelian elements. (In fact, it is also the closed linear span of the abelian elements.) Again, one can recursively define the ideals $I_{\alpha}\left(A\right)$ for $\alpha\leq\omega_{1}$, and again $A$ is type $I$ if and only if $A=I_{\alpha}\left(A\right)$ for some $\alpha<\omega_{1}$ [7, IV.1.1.12]. Furthermore, $I_{\omega_{1}}\left(A\right)=J_{\omega_{1}}\left(A\right)$ is the largest type $I$ C*-subalgebra of $A$ [7, IV.1.1.12 and IV.1.3.9].
@@
see [7, IV.1.4.8]. The element $x$ has continuous trace if $\hat{x}$ takes values in $[0,n]$ for some $n\in\mathbb{N}$ and it is continuous. Then the set $\mathfrak{m}\left(A\right)_{+}$ of positive continuous trace elements of $A$ is the positive part of an ideal $\mathfrak{m}\left(A\right)$ [7, IV.1.4.11 ]. A C*-algebra $A$ has continuous trace if $\mathfrak{m}\left(A\right)$ is dense in $A$.
@@
The equivalence of (1)—(7) is the content of [7, IV.1.5.7, IV.1.5.12] and [48, Theorem 6.9.7]; see also [7, IV.1.5.7]. The equivalence of (a)—(c) is the content of [28, Theorem 4]. ∎
@@
In the unital case, a liminary C*-algebra is necessarily preliminary [7, Section IV.1.3]. If $A$ is any preliminary nontrivial separable C*-algebra, then $A\otimes K\left(H\right)$ is liminary and not preliminary.

### v4

@@ 1. Introduction
This paper is a contribution to the study of the structure and classification of separable C*-algebra. We focus on the class of so-called liminary C*-algebras, also known as liminal or CCR. These are precisely the C*-algebras whose irreducible representations on a Hilbert space have the property that their image coincides with the algebra of compact operators. The more generous requirement that the image contains the algebra of compact operators yields the notion of postliminary C*-algebra, also known as postliminal or Type I. Within the class of postliminary C*-algebras, liminary C*-algebras are precisely those whose primitive spectrum endowed with the Jacobson topology has closed points—i.e., every prime ideal is maximal within the class of prime (or, equivalently, all) ideals [31, Corollary 2.6].
The study of liminary and postliminary C*-algebras goes back to the early days of C*-algebra theory. Its forefathers are Kaplansky [42, 41] and Mackey [52, 49, 50, 51], who pioneered the study of liminary and postliminary C*-algebras in the 1950s. Their motivation included the classification problem for C*-algebras and their irreducible representations. The classification problem of irreducible representations of locally compact Hausdorff topological groups can be seen as a particular instance of the one for C*-algebras, by considering the corresponding group C*-algebra. The theory of liminary and postliminary C*-algebras was further developed in the 1960s by Dixmier [16, 17, 15, 19], Effros [22, 23], Fell [26, 28, 27], and Glimm [32] among others.
An upshot of these works is Glimm’s Theorem from [32] characterizing separable postliminary C*-algebras in terms of the corresponding classification problem for irreducible representations: a separable C*-algebra is postliminary if and only if its irreducible representations are concretely classifiable up to unitary equivalence, and the corresponding Borel structure induced on the spectrum of the C*-algebra is standard.
Since then, the study of the representation theory of C*-algebras has divided into two streams, divided by the type $I$ versus non-type $I$ dichotomy. In the non-type $I$ case, the study of the classification problem for irreducible C*-algebras and their “definable spectra” has refined the analysis from the perspective of Borel complexity theory [44, 68, 25, 35]. In the type $I$ case, more stringent notions have been introduced in the attempt to obtain a complete description of the C*-algebras under considerations and their spectra.
@@
As observed in [24], a type $I_{0}$ and, more generally, a type $I_{n}$ C*-algebra is necessarily liminary, and the same holds for arbitrary countable ordinals. It is also remarked in [24] that there exist liminary C*-algebras that are not type $I_{n}$ for any $n<\omega$. We prove that by considering arbitrary countable ordinals one obtains a hierarchy that is complete, and includes all liminary C*-algebras.
@@
We also prove that the hierarchy is strict, namely, for any countable ordinal $\alpha$ there exist separable liminary C*-algebras that are type $I_{\alpha}$ and not type $I_{\beta}$ for any $\beta<\alpha$. Such C*-algebras can be chosen to be scattered (particularly, approximately finite-dimensional) with a countable compact metrizable spectrum, and no infinite-dimensional irreducible representations. Other examples of such C*-algebras were previously considered by Lazar and Taylor [47].
@@
We extend the Enders–Shulman characterization to type $I_{\alpha}$ C*-algebras for an arbitrary countable ordinal $\alpha$. This is obtained in terms of the (ordinal-valued) Pedersen rank of an element of a C*-algebra, generalizing the notion of abelian element (corresponding to the rank 0 case) and global rank $n$ (corresponding to the case of finite ordinals). In turn, this yields a corresponding notion of $\alpha$-subhomogeneous C*-algebra for an arbitrary countable ordinal $\alpha$. Whereas $n$-subhomogeneous C*-algebras have all irreducible representations of dimension at most $n$, an $\alpha$-subhomogeneous C*-algebra can have irreducible representations of arbitrary large, albeit finite, dimension when $\alpha$ is infinite.
We define a C*-algebra to be preliminary if it has no infinite-dimensional representations. A preliminary C*-algebra is, in particular, liminary. This class of C*-algebras has received attention recently, in the work of Courtney and Shulman [12]; see also [57, 56, 63, 66, 46, 38, 39, 67, 34, 10]. As in the case of type $I_{\alpha}$ C*-algebras within liminary C*-algebras, we prove that the hierarchy of $\alpha$-subhomogeneous C*-algebras within preliminary C*-algebras is complete.
@@
Again, this hierarchy is strict, as for any countable ordinal $\alpha$ one can find $\alpha$-subhomogeneous C*-algebras that are not $\beta$-subhomogeneous for any $\beta<\alpha$. Furthermore, one can even find such examples to be scattered with countable compact metrizable spectrum.
As a reference for notions pertaining to C*-algebras we refer the reader to the monographs [9, 60, 20]. We will also invoke results from general topology and descriptive set theory, as can be found in the manuals [43, 30].
The rest of this paper is divided into eleven sections, apart from this introduction. In Section 2 we recall some fundamental notions concerning the classes of C*-algebras that we will consider. In Section 3 we will introduce some notions from topology and descriptive set theory. In Section 4 we present the Fell compactification construction, and record some of its properties. The notion of weighted space is introduced in Section 5, whereas Section 6 observes that the spectra of preliminary C*-algebras are weighted spaces. Section 7 introduces the notions of Fell and Pedersen rank for C*-algebras, and the corresponding notions of type $I_{\alpha}$ and $\alpha$-subhomogeneous C*-algebra for an arbitrary countable ordinal $\alpha$. Section 8 extends their characterization from finite to countable ordinals. The fact that the hierarchies of type $I_{\alpha}$ and $\alpha$-subhomogeneous C*-algebras are complete is proved in Section 9. In he Section 10 we recall terminology pertaining to countable trees, to be used in the final section. Finally, Section 11 produces examples of preliminary C*-algebras of arbitrarily high Pedersen and Fell rank.
@@ Acknowledgments
We are thankful to Jeffrey Bergfalk, Luigi Caputi, Alessandro Codenotti, Ivan Di Liberti, Eusebio Gardella, Ilja Gogic, Alexander Kechris, Aristotelis Panagiotopoulos, Ebrahim Samei, Tatiana Shulman, and Joseph Zielinski for several helpful conversations.
@@ 2. Some C*-algebra notions
In this section we recall some notions from C*-algebra theory, and particularly the notions of liminary and postliminary C*-algebra, as well as of scattered C*-algebra; see also [60, 20, 9].
@@ 2.1. Spectrum of a C*-algebra
Let $A$ be a separable C*-algebra. The quasi-state space $Q\left(A\right)$ of $A$ is the space of positive linear functionals on $A$ of norm at most $1$, which is a w*-closed subspace of the dual of $A$ [60, Section 3.2.1]. The space $P\left(A\right)$ of non-zero extreme points of $Q\left(A\right)$ is a $G_{\delta}$ subspace of $Q\left(A\right)$, whence Polish when endowed with the subspace topology [60, Proposition 4.3.2]. The elements of $P\left(A\right)$ are called pure states of $A$.
Let $A$ be a separable C*-algebra. The Borel T-structure (T stands for topological) on $\hat{A}$ is the $\sigma$-algebra generated by the open sets in the Jacobson topology, while the Borel M-structure (M stands for Mackey) on $\hat{A}$ is the one induced from the standard Borel structure on $P\left(A\right)$ by the map $P\left(A\right)\rightarrow\hat{A}$ mapping a pure state $\phi$ to the kernel of the corresponding irreducible representation $\pi_{\phi}$ [60, Theorem 4.3.3]. Then we have that the Borel T-structure is weaker than the Borel M-structure, but they coincide (and they are standard) when $A$ is type $I$ [60, Proposition 6.3.2].
Suppose that $x\in A$. Then $x$ defines a lower semi-continuous function $\check{x}:\mathrm{Prim}\left(A\right)\rightarrow\mathbb{R}$ by $\check{x}\left(J\right):=\left\|x+J\right\|_{A/J}$. In fact, the Jacobson topology on $\mathrm{Prim}\left(A\right)$ is the weakest topology that makes $\check{x}$ lower semi-continuous for all $x$ in (some countable dense subset of) $A$; see [60, Section 4.4].
If $A$ and $B$ are separable C*-algebras, and at least one between $A$ and $B$ is type $I$, then there exists a canonical continuous map from $\hat{A}\times\hat{B}$ to the spectrum of $A\otimes B$ [9, IV.3.4.22]. This is in fact a homeomorphism [9, IV.3.4.28], and it induces a homeomorphism
@@
In particular, $\mathrm{Prim}\left(A\right)$ and $\mathrm{Prim}\left(A\otimes K\left(H\right)\right)$ are homeomorphic, since $\mathrm{Prim}\left(K\left(H\right)\right)$ is a singleton [9, IV.1.2.2].
@@ 2.2. Liminary and postliminary C*-algebras
Let $A$ be a separable C*-algebra, and $\pi$ be an irreducible representation of $A$ on a Hilbert space $H$. Then $\pi$ is CCR if $\pi\left(A\right)\subseteq K\left(H\right)$ (which implies $\pi\left(A\right)=K\left(H\right)$), and GCR if $\pi\left(A\right)\cap K\left(H\right)\neq 0$ (and hence $K\left(H\right)\subseteq\pi\left(A\right)$); see [9, IV.1.3.1]. A C*-algebra is CCR (respectively, GCR) if every irreducible representation of $A$ is CCR (respectively, GCR). A C*-algebra is liminary if and only if it is CCR; see [9, IV.1.3.1]. Recall also that an element $x$ of $A$ is abelian if the hereditary subalgebra $\overline{xAx^{\ast}}$ it generates is commutative [9, IV.1.1.1]. A Fell C*-algebra $A$ (also called type $I_{0}$ C*-algebra) is a C*-algebra generated by its abelian elements [9, IV.1.1.6]. Every Fell C*-algebra is liminary [9, IV.1.3.2]. A separable C*-algebra is elementary if it is isomorphic to $K\left(H\right)$ for some separable Hilbert space $H$ [9, IV.1.2.1]. Liminary C*-algebras were originally introduced by Kaplansky in [41]; see [60, Section 6.2.13]. Fell C*-algebras were introduced by Pedersen [60, Section 6.1.14], and can be characterized in terms of Fell’s condition [9, IV.1.4.17].
@@
Let $A$ be a separable C*-algebra and $\mathcal{C}$ be a class of C*-algebras. A decomposition series for $A$ with subquotients in $\mathcal{C}$ is a chain $\left(R_{\alpha}\right)_{\alpha<\omega_{1}}$ of closed ideals of $A$ such that:
@@
If $A$ is a separable C*-algebra, then it has a largest liminary closed ideal (i.e., a largest element in the collection of closed ideals that are liminal as C*-algebras); see [9, IV.1.3.9]. This can be explicitly defined as the ideal $J_{1}\left(A\right)$ comprising the $x\in A$ such that for every irreducible representation $\pi$ of $A$, $\pi\left(x\right)$ is a compact operator. Then one can recursively define a chain $\left(J_{\alpha}\left(A\right)\right)_{\alpha\leq\omega_{1}}$ of closed ideals, continuous at limits, such that
@@
for $\alpha\leq\beta$. The C*-algebra $A$ is postliminary if $A=J_{\alpha}\left(A\right)$ for some $\alpha<\omega_{1}$; see [9, IV.1.3.9].
A separable C*-algebra also has a largest Fell ideal [9, IV.1.1.8]. This can be explicitly defined as the C*-subalgebra $I_{1}\left(A\right)$ of $A$ generated by its abelian elements. (In fact, it is also the closed linear span of the abelian elements.) Again, one can recursively define the ideals $I_{\alpha}\left(A\right)$ for $\alpha\leq\omega_{1}$, and again $A$ is type $I$ if and only if $A=I_{\alpha}\left(A\right)$ for some $\alpha<\omega_{1}$ [9, IV.1.1.12]. Furthermore, $I_{\omega_{1}}\left(A\right)=J_{\omega_{1}}\left(A\right)$ is the largest type $I$ C*-subalgebra of $A$ [9, IV.1.1.12 and IV.1.3.9].
@@
see [9, IV.1.4.8]. The element $x$ has continuous trace if $\hat{x}$ takes values in $[0,n]$ for some $n\in\mathbb{N}$ and it is continuous. Then the set $\mathfrak{m}\left(A\right)_{+}$ of positive continuous trace elements of $A$ is the positive part of an ideal $\mathfrak{m}\left(A\right)$ [9, IV.1.4.11 ]. A C*-algebra $A$ has continuous trace if $\mathfrak{m}\left(A\right)$ is dense in $A$.
@@
The equivalence of (1)—(7) is the content of [9, IV.1.5.7, IV.1.5.12] and [60, Theorem 6.9.7]; see also [9, IV.1.5.7]. The equivalence of (a)—(c) is the content of [32, Theorem 4]. ∎
@@
In the unital case, a liminary C*-algebra is necessarily preliminary [9, Section IV.1.3]. If $A$ is any preliminary nontrivial separable C*-algebra, then $A\otimes K\left(H\right)$ is liminary and not preliminary.
2.3. Scattered C*-algebras
The class of scattered C*-algebras has originally been introduced by Jensen [36, 37]. A separable C*-algebra $A$ is scattered if the dual space $\hat{A}$ is a scattered topological space; see [37, Corollary 3]. As in the commutative case, separable scattered C*-algebras admit numerous equivalent characterization.
Proposition 2.4.
  1. (1)
    AA is scattered, i.e., the dual space $\hat{A}$ is scattered;
  2. (2)
    AA admits a decomposition series $\left(E_{\alpha}\right)_{\alpha<\omega_{1}}$ with elementary subquotients;
  3. (3)
    AA is a type $I$ AF C*-algebra and $\mathrm{Prim}\left(A\right)$ is scattered;
  4. (4)
    AA is a type $I$ C*-algebra and $\mathrm{Prim}\left(A\right)$ is countable;
  5. (5)
    the dual space $\hat{A}$ is countable.
Proof.
[…]
Lemma 2.5.
Let $A$ be a separable C*-algebra, $B\subseteq A$ be a C*-subalgebra, and $J\subseteq A$ be a closed ideal.
  1. (1)
    If $A$ is liminal (respectively, postliminal), then so are $B$ and $A/J$;
  2. (2)
    If $A$ and $A/J$ are postliminal, then so is $A$;
  3. (3)
    AA is scattered if and only if $J$ and $A/J$ are scattered.
Proof.
(1) See [9, IV.1.3.10] for the conclusion about $B$. The one about $A/J$ is obvious.
(2) This follows from the fact that an irreducible representation of $A$ is either zero or irreducible when restricted to $J$ [9, II.6.1.6].
(3) This is the content of [36, Proposition 2.4]. ∎
If $X$ is a scattered locally compact Polish space, then the C*-algebra $C_{0}\left(X\right)$ of continuous complex-valued functions on $X$ vanishing at infinity is a commutative scattered separable C*-algebra with (primitive) spectrum $X$. Conversely, the spectrum $X$ of a commutative scattered separable C*-algebra $A$ is a scattered locally compact Polish space with $C_{0}\left(X\right)\cong A$. Thus, arbitrary (noncommutative) separable C*-algebras can be seen as the “noncommutative generalization” of scattered locally compact Polish spaces.